OpenStudy (anonymous):

Help Please! <3 Tell whether or not the given point is a solution to the system of equations. (4,2) y=0.5+1 y=0.6+0.6

jimthompson5910 (jim_thompson5910):

x is missing in both equations

OpenStudy (anonymous):

Sorry its y=0.5x+1 y=0.6x+0.6

jimthompson5910 (jim_thompson5910):

y = 0.5x + 1 2 = 0.5*(4) + 1 ... plug in x = 4 and y = 2 now simplify both sides and tell me what you get

OpenStudy (jim766):

you can solve this way y = .5x + 1 y = .6x +.6 y = y .5x +1 = .6x + .6 solve for x then sub to find y

OpenStudy (anonymous):